# On the reconstruction limits of complex networks

### v1

@@ Abstract
Network reconstruction consists in retrieving the—hidden—interaction structure of a system from empirical observations such as time series. Many reconstruction algorithms have been proposed, although less research has been devoted to describe their theoretical limitations. To this end, we adopt an information-theoretical point of view and define the reconstructability—the fraction of structural information recoverable from data. The reconstructability depends on the true data generating model which is shown to set the reconstruction limit, i.e., the performance upper bound for all algorithms. We show that the reconstructability is related to various performance measures, such as the probability of error and the Jaccard similarity. In an empirical context where the true data generating model is unknown, we introduce the reconstruction index as an approximation of the reconstructability. We find that performing model selection is crucial for the validity of the reconstruction index as a proxy of the reconstructability, and illustrate how it assesses the reconstruction limit of empirical time series and networks.
@@ I Introduction
Complex systems such as the brain are naturally represented by complex networks, which encapsulate intricate interactions between neurons or brain regions Bassett and Sporns (2017); Bassett et al. (2018); Lynn and Bassett (2019); Sporns (2013). Network representation unlocks a variety of tools with the potential to unravel not only brain functions and diseases Supekar et al. (2008); Hlinka and Coombes (2012); Forrester et al. (2020), but also gene expressions Wang et al. (2006), epidemics Peixoto (2019); Prasse et al. (2020) and financial distress Musmeci et al. (2013). The problem is that such network representations are seldom measurable experimentally. For example, the collected data are often indirect observations of the interactions, taking the form of counts of interactions or times series. Moreover, these data are noisy, thereby making the network reconstruction task even more intricate Newman (2018); Peixoto (2018); Young et al. (2020); Peel et al. (2022).
@@
Another promising avenue for network reconstruction involves quantifying uncertainty in real-world networks. In this setting, noisy pairwise observations are used to predict missing edges Clauset et al. (2008); Guimerà and Sales-Pardo (2009); Lü and Zhou (2011), estimate the edge uncertainty Newman (2018) and reconstruct the network altogether Peel et al. (2022). As for network reconstruction from time series, heuristics have also been considered for pairwise data (for example, in Ref. Goldberg and Roth (2003)). Recently, there has also been a resurgence in the interest towards Bayesian frameworks. For instance, Ref. Young et al. (2020) proposed a general and scalable Bayesian procedure to infer networks leveraging the conditional independence of the edges, which was then applied to a plant-pollinator network Young et al. (2021). Reference Lizotte et al. (2023) extended this framework to the reconstruction of hypergraphs with noisy observations and showed the benefit of including higher-order interactions for modeling pairwise measurements. Other works used the modular structure of complex networks to improve the performance of their models Clauset et al. (2008); Guimerà and Sales-Pardo (2009); Peixoto (2018). To this date, the field of reconstruction of noisy networks remains a flourishing one.
@@
As more technical progress is being made, more work is being dedicated to the theoretical challenges of network reconstruction. For instance, Ref. Peel et al. (2022) proposed a unifying framework for linking network data to network science theories, in which Bayesian network reconstruction is core and where the suitability of the models, they argue, is essential for network reconstruction. Additionally, Ref. Prasse and Mieghem (2022) found that network reconstruction, on the basis of predicting the outcome of a deterministic dynamical process, can lead to a wide range of networks. The relationship between reconstructability and predictability in networked processes, and in turn their duality, has also been corroborated in Ref. Murphy et al. (2024) using information-theoretic tools. These theoretical findings highlight different facets of the same phenomenon: the limit of network reconstructability. To what extent is a network reconstructible, and how descriptive are the networks we observe in—and reconstruct from—nature? Answering these questions fundamentally changes how empirical networks are analyzed, i.e., by acknowledging their intrinsic uncertainty.
@@
In this work, we investigate the very question of network reconstructability. In Sec. II, we describe more precisely the network reconstruction problem using a Bayesian perspective. We identify the crucial role that data generation plays in our ability to reconstruct networks, and link it to the reconstruction limit. Section III formalizes this limit in information-theoretic terms. We find that the mutual information, between the graph and the data used to reconstruct, is the limiting factor of network reconstruction. Section IV generalizes the mutual information to a realistic, data-driven scenario—where only the observations are known. As in Ref. Peel et al. (2022), we find that the quality of the model to describe the observations is critical for our ability to describe the reconstruction limit. Finally, in Sec. V we apply our framework to time series and empirical networks.
@@
Fig. 1: Illustration of the network reconstruction context from (a, b) a theoretical perspective and (c, d) an empirical perspective. Panel (a) sketches how the true data generating model (TDG) $M^{*}$ operates, first by generating a graph, then by encoding it into the observations, and finally using these to decode—or reconstruct—the graph. The thickness of the contour line around each graph and data example indicates the probabilities $P(G^{*})$ (top and bottom layers) and $P(X^{*})$ (middle layer). The thickness of the edges connecting the graphs to the data illustrate the likelihood of the TDG $P(X^{*}|G^{*})$, and those connecting the data to a reconstructed graphs, some posterior distribution $P(G|X=X^{*})$. In panel (b), we illustrate in red the reconstructible information, utilizing an information-theoretic perspective. This information is part of the total information of $G^{*}$ and $X^{*}$—in blue and orange, respectively—and is also a fraction of the partial information of $G^{*}$ needed to completely reconstruct it (blue and red). Panels (c, d) show the analog of (a, b) when model $M^{*}$ is unknown, where in panel (c) a single datum is accessible and reconstruction is done by a candidate model $M$, a priori different from $M^{*}$. In panel (d), we illustrate how $M$ and $M^{*}$ may overlap in the information they reconstruct—the information intersection (i.e., the correctly recovered information) and difference (i.e., the missing or spurious information). The reconstructability $U(G^{*}|X^{*})$ and the reconstruction index $\Psi_{M}$ are defined in subsection III.2 and subsection IV.1, respectively.
II Network reconstruction
@@
We formulate the network reconstruction problem as follows. Let $g^{*}$ be some graph of $N$ nodes, belonging to a set $\mathcal{G}$ that represents the structure of the interactions between each pair of components in a system. The graph may be directed and weighted Peixoto (2024), but we restrict our discussion to undirected and unweighted, for simplicity. This graph structure is a priori unknown to us, although it is indirectly observed through some data, denoted $x^{*}$, which may take any value in the set $\mathcal{X}$. This data can take many forms—time series, pairwise measurements, etc.—and we assume it to be generated using $g^{*}$. The goal of network reconstruction is to infer the graph $g^{*}$ from the data $x^{*}$.
@@
Taking a Bayesian perspective, the plausibility of a given graph $g\in\mathcal{G}$ is given by the posterior probability $P(G=g|X)$, i.e., the output of the Bayesian inference procedure. The reconstruction model describes two discrete random variables $G$ and $X$ that represent the graphs and the data respectively, and thus defines their joint probability mass function. By virtue of Bayes’ theorem, the posterior $P(G|X)$ is factored as a product of the graph prior $P(G)$ and data likelihood $P(X|G)$:
@@
II.1 Data generation process
@@
In Bayesian modeling, a reconstruction model, composed of two random variables $G$ and $X$, reflects our assumptions about how the unobserved graph and observed data came to be. In other words, the model $M=(G,X)$ represents a generative process for the pairs $(g^{*},x^{*})$. Accordingly, there are many reconstruction models that may explain the data to various degrees of correctness. The true data generating (TDG) model is the generative process $M^{*}=(G^{*},X^{*})$ which truly produced the graph $g^{*}$ and observed data $x$ with probabilities $P(G^{*}=g^{*})$ and $P(X^{*}=x|G^{*}=g^{*})$, respectively. While the random variables of the TDG and of the reconstruction models are independent, they are related in the complete reconstruction process through the following joint probability:
$\begin{split}P(G^{*}=g^{*},X^{*}=x^{*},G=g&|X=x^{*})\\ =P(G^{*}=g^{*})&P(X^{*}=x^{*}|G=g^{*})\\ \times&P(G=g|X=x^{*})\,.\end{split}$ (2)
Intuitively, this expression states that, while $(g^{*},x^{*})$ are produced by $M^{*}$, $M$ may reconstruct any graph $g$ under the condition that its data variable is set to $x^{*}$. In general, we assume a process such as $M^{*}$ exists. Additionally, the model $M$ may be equal in distribution to $M^{*}$—which amounts to performing reconstruction with the true posterior distribution $P(G^{*}|X^{*})$—, although it will not generally be the case.
From an information-theoretic perspective, data generation encodes information about the graph $G^{*}$ into—potentially noisy—observations $X^{*}$, while network reconstruction decodes these observations back into a graph $G$, as shown in Fig. 1(a–b). The encoding of $G^{*}$ into $X^{*}$ is generally lossy, meaning that only a fraction of its information can be recovered; the rest being lost in the process. In turn, any reconstruction model $M$ different in distribution from $M^{*}$ may therefore recover a fraction of the reconstructible information, while potentially introducing spurious information through their inductive biases [see Fig. 1(c–d)], resulting in a degradation of performance. This is well shown in Fig. 2 through reconstruction performance, where we compare the area under the receiver operating characteristic curve (AUC) of the TDG reconstruction model—i.e., whose distribution is given by that of $M^{*}$—, with the AUC of three heuristic reconstruction algorithms. In this example, the TDG model is a simple Erdős-Rényi (ER) graph, whose graph prior is
@@
with $E$ being the (given) number of edges in the graph (see Appendix A for details). Then, the data consists in time series of binary node activities, generated using the Glauber dynamics model Glauber (1963) in panel (a) and the Susceptible-Infected-Susceptible (SIS) dynamics model Pastor-Satorras et al. (2015) in panel (b) (see Appendix B for details). As a comparison, we use three different well-known reconstruction algorithms: the correlation matrix method Kramer et al. (2009), Granger causality method Schreiber (2000) and the transfer entropy method Seth (2005) (see Appendix C for details). The results in Fig. 2 show quite unambiguously and unsurprisingly that the TDG model outperforms the reconstruction heuristics.
Yet, even the TDG reconstruction model cannot reconstruct the graph perfectly. For instance, in Fig 2(b), the AUC of the TDG model tends to $\frac{1}{2}$—equivalent to random guessing—when the coupling also goes to zero. In this scenario, $X^{*}$ and $G^{*}$ are independent and it is actually impossible to reconstruct the graph, since any graph could have generated the data with the exact same probability. The same phenomenon occurs to a lesser extent for the other coupling values as well as for the SIS dynamics, where the TDG model performance is imperfect for every infection probabilities. These imperfections are attributed to the lost information in the encoding of $G^{*}$; no model can extract more information than what is contained in $X$. In practice, the encoding’s loss stems from many sources, for example noise in the dynamics and degeneracy, where many networks lead to similar dynamics. The degeneracy phenomenon is well-established in computational neuroscience Prinz et al. (2004); Cropper et al. (2016) and has more recently appeared in network science Prasse and Mieghem (2022) too. A reconstruction limit independent of the reconstruction algorithm clearly exists, where a perfect reconstruction is simply not attainable even in the best-case scenario. This is a key insight that we will explore in the following sections.
II.2 Reconstructing a single edge
To gain better intuition about this reconstruction limit, we consider the reconstruction of a graph composed of a single edge. Let $G^{*}$ be a random graph composed of two nodes, that may be connected by a single edge with probability $p$, and disconnected with probability $1-p$. This edge is observed through a noisy process $X^{*}=(X_{1},...,X_{T})$ with $T$ time steps, where $X_{i}$ is a binary variable that takes the value $1$ if the edge has been observed and 00 otherwise. We assume that the noisy process can induce true positives and false positives, each with known probabilities $q$ and $r$, respectively, making the reconstruction problem more challenging. The likelihood $P(n|a)$ that the edge has been observed $n$ times, given that it is present ($a=1$) or not ($a=0$), is a binomial distribution:
@@ II.2 Reconstructing a single edge
Note that this model possesses a symmetry where interchanging $q$ and $r$ leaves the likelihood unchanged if we map $a\to 1-a$ and vice-versa. However, we avoid this non-identifiability issue by not inferring $p$ and $r$.
@@
Figure 3 shows the behavior of $P(a=1|n)$ when varying the number $n$ of times the edge is observed and the true positive probability $q$. Assuming that $r<q$, observed edges are mostly true positives and thus the edge is predicted to exist if $n$ is sufficiently large; otherwise, it is not—i.e., the expected number of true and false positives don’t match the observations. Conversely, if $r>q$, then most observed edges are false positives, meaning that $g^{*}$ is more likely to contain an edge when $n$ is small. Interestingly, the edge becomes more challenging to reconstruct as $q$ gets closer to $r$, where the probability to reconstruct the edge approaches $\frac{1}{2}$ (see Fig. 3(a)). In this regime, $a$ and $1-a$ are interchangeable and it becomes impossible to tell if the edge exists or not—any attempt at reconstructing this graph would be unfruitful. This is precisely the intuition we want to capture with the reconstruction limit; When is there enough information to properly reconstruct the structure, or to what extent is a system’s structure reconstructible? In the next section, we present an information-theoretic framework that quantifies this limit.
III Information-theoretic reconstruction limits
As discussed earlier, we can think of the TDG process $X^{*}$ as a noisy encoding of the true graph $G^{*}$, which encrypts information about it. This amount of encrypted information is fundamentally limiting our ability to reconstruct $G^{*}$ accurately; it is impossible to recover more information than what is contained in the data. This also means that the limit is independent of the reconstruction models or algorithms. Any reconstruction algorithm therefore attempts to extract as much of the encrypted information as possible, some being more efficient than others.
III.1 Entropy

### v2

@@ Abstract
Network reconstruction consists in retrieving the hidden interaction structure of a system from observations. Many reconstruction algorithms have been proposed, although less research has been devoted to describe their theoretical limitations. In this work, we adopt an information-theoretic perspective and define the reconstructability: The fraction of structural information recoverable from data. The reconstructability depends on the true data generating (TDG) model which is shown to set the reconstruction limit: any algorithm can perform, on average, at best like the TDG model. We show that the reconstructability is related to various performance measures, such as the probability of error and the Jaccard similarity. In an empirical context where the TDG model is unknown, we introduce the reconstruction index as an approximation of the reconstructability. We find that performing model selection is crucial for the validity of the reconstruction index as a proxy of the reconstructability of empirical time series and networks.
@@ I Introduction
Complex systems, such as the brain, are naturally represented by complex networks that encapsulate intricate interactions between neurons or brain regions Bassett and Sporns (2017); Bassett et al. (2018); Lynn and Bassett (2019); Sporns (2013). Network representation unlocks a variety of tools with the potential to unravel not only brain functions and diseases Supekar et al. (2008); Hlinka and Coombes (2012); Forrester et al. (2020), but also gene expressions Wang et al. (2006), epidemics Peixoto (2019); Prasse et al. (2020) and financial distress Musmeci et al. (2013). The problem is that such network representations are seldom measurable experimentally. For example, the collected data are often indirect observations of the interactions, taking the form of counts of interactions or times series. Moreover, these data are noisy, thereby making the network reconstruction task even more intricate Newman (2018); Peixoto (2018); Young et al. (2020); Peel et al. (2022).
@@
Another promising avenue for network reconstruction involves using pairwise observations for quantifying the uncertainty of empirical graphs. In this setting, noisy pairwise observations are used to predict missing edges Clauset et al. (2008); Guimerà and Sales-Pardo (2009); Lü and Zhou (2011), estimate the edge uncertainty Newman (2018) and reconstruct the network altogether Peel et al. (2022). As for network reconstruction from time series, heuristics have also been considered for pairwise data (for example, in Ref. Goldberg and Roth (2003)). Recently, there has also been a resurgence in the interest towards Bayesian frameworks. For instance, Ref. Young et al. (2020) proposed a general and Bayesian procedure to infer networks leveraging the conditional independence of the edges, which was then applied to a plant-pollinator network Young et al. (2021). Reference Lizotte et al. (2023) extended this framework to the reconstruction of hypergraphs with noisy observations and showed the benefit of including higher-order interactions for modeling pairwise measurements. Other works used the modular structure of complex networks to improve the performance of their models Clauset et al. (2008); Guimerà and Sales-Pardo (2009); Peixoto (2018). To this date, the field of reconstruction of noisy networks remains a flourishing one.
@@
As more technical progress is being made, more work is being dedicated to the theoretical challenges of network reconstruction. For instance, Ref. Peel et al. (2022) proposed a unifying framework for linking network data to network science theories, in which Bayesian network reconstruction is core and where they argue the suitability of the models is essential for network reconstruction. Additionally, Ref. Prasse and Mieghem (2022) found that network reconstruction, on the basis of predicting the outcome of a deterministic dynamical process, can lead to a wide range of networks. This aligns with the observations of Ref. Angulo et al. (2017) and earlier computational neuroscience findings Prinz et al. (2004) of network degeneracy Cropper et al. (2016), where diverse synaptic connection patterns can yield similar neuronal activity, illustrating the non-unique relationship between network structure and function.
The structure-function relationship has recently been investigated in a broader context for random networked processes Murphy et al. (2024) using information-theoretic tools. The study uncovered a duality between reconstructability and predictability, showing that in certain parameter ranges, an increase in predictability corresponds to a decrease in reconstructability, and vice versa. These theoretical findings provide a unifying framework for addressing two a priori unrelated questions : To what extent is a network reconstructible, and how descriptive are the networks we observe in—and reconstruct from—nature?
@@
In this work, we focus on the question of the reconstructability of networks. In Sec. II.1, we describe more precisely the network reconstruction problem using a Bayesian perspective. We identify the crucial role that data generation plays in our ability to reconstruct networks, and link it to the reconstruction limit. Section II.2 formalizes this limit in information-theoretic terms. We find that the mutual information, between the graph and the data used to reconstruct, is the limiting factor of network reconstruction. Section II.3 generalizes the mutual information to a realistic, data-driven scenario—where only the observations are known. As in Ref. Peel et al. (2022), we find that the quality of the model to describe the observations is critical for our ability to describe the reconstruction limit. Finally, in Sec. II.4 we apply our framework to time series and empirical networks.
@@
Fig. 1: Illustration of the network reconstruction context from (a, b) a theoretical perspective and (c, d) an empirical perspective. Panel (a) sketches how the true data generating model (TDG) $M^{*}$ operates, first by generating a graph, then by encoding it into the observations, and finally using these to decode—or reconstruct—the graph. The thickness of the contour line around each graph and data example indicates the probabilities $P(G^{*})$ (top and bottom layers) and $P(X^{*})$ (middle layer). The thickness of the edges connecting the graphs to the data illustrate the likelihood of the TDG $P(X^{*}|G^{*})$, and those connecting the data to a reconstructed graphs, some distribution $P(\hat{G}|X^{*})$. In panel (b), we illustrate in red the reconstructible information, utilizing an information-theoretic perspective. This information is part of the total information of $G^{*}$ and $X^{*}$—in blue and orange, respectively—and is also a fraction of the partial information of $G^{*}$ needed to completely reconstruct it (blue and red). Panels (c, d) show the analog of (a, b) when model $M^{*}$ is unknown, where in panel (c) a single datum is accessible and reconstruction is done by a candidate model $M$, a priori different from $M^{*}$. In panel (d), we illustrate how $M$ and $M^{*}$ may overlap in the information they reconstruct—the information intersection (i.e., the correctly recovered information) and difference (i.e., the missing or spurious information). The reconstructability $U(G^{*}|X^{*})$ and the reconstruction index $\Psi_{M}$ are defined in subsection II.2.2 and subsection II.3.1, respectively.
II Results
II.1 Network reconstruction
@@
We formulate the network reconstruction problem as follows. Let $g^{*}\in\mathcal{G}$ be some graph of $N$ nodes that represents the structure of the interactions between each pair of components in a system, where $\mathcal{G}$ is the set of all graphs of $N$ nodes. The graph may be directed and weighted Peixoto (2024), but we restrict our discussion to undirected and unweighted, for simplicity. This graph structure is a priori unknown to us, although it is indirectly observed through some data, denoted $x^{*}$, which may take any value in the set $\mathcal{X}$. This data can take many forms—time series, pairwise measurements, etc.—and we assume it to be generated using $g^{*}$. The goal of network reconstruction is to infer the graph $g^{*}$ from the data $x^{*}$.
@@
Taking a Bayesian perspective, the plausibility of a given graph $g\in\mathcal{G}$, given the observations $x^{*}$, is described by the posterior probability $P(G=g|X=x^{*})$, i.e., the output of the Bayesian inference procedure. A Bayesian reconstruction model is a generative process that consists of two discrete random variables $G$ and $X$, representing the graphs and the data respectively, and thus defines their joint probability mass function $P(G,X)=P(G)P(X|G)$, where $P(G)$ is the graph prior and $P(X|G)$, the data likelihood. By virtue of Bayes’ theorem, the posterior $P(G|X)$ is factored as follows:
@@
II.1.1 Data generation process
@@
A Bayesian reconstruction model, composed of the two random variables $G$ and $X$, reflects our assumptions about how the unobserved graph and observed data came to be. In other words, the model $M=(G,X)$ represents a generative process for the pairs $(g^{*},x^{*})$. Accordingly, there are many reconstruction models that may describe the data to various degrees of correctness. Throughout this work, we assume the existence of a unique generative process, referred to as the true data-generating (TDG) model $M^{*}=(G^{*},X^{*})$, which truly produced the graph $g^{*}$ and the observed data $x^{*}$ with probabilities $P(G^{*}=g^{*})$ and $P(X^{*}=x^{*}|G^{*}=g^{*})$, respectively. In turn, any reconstruction model may be described by a reconstructed random graph $\hat{G}$, that depends on $X^{*}$. The complete process consisting of the graph and data generation followed by the reconstruction of the graph is therefore described by the random variable triplet $(G^{*},X^{*},\hat{G})$, whose joint probability distribution is
$P(G^{*},X^{*},\hat{G})=P(G^{*})P(X^{*}|G^{*})P(\hat{G}|X^{*})\,.$ (2)
In general, the distribution $P(\hat{G}|X^{*})$ may be any distribution over $\mathcal{G}$, but for Bayesian models such as $M$, it is precisely given by the posterior of $M$:
$P(\hat{G}=g|X^{*}=x^{*})=P(G=g|X=x^{*})\,$ (3)
for all $g\in\mathcal{G}$, such that $P(G|X)$ is given by Eq. (1). Note that the reconstructed random graph $\hat{G}$ and the random graph $G$ of model $M$ conceptually describe two different quantities, although they are related through Eq. (3). Indeed, $\hat{G}$ appears in the reconstruction process involving the TDG and $G$ is part of a completely separate generative process. In other words, $\hat{G}$ depends explicitly on $M^{*}$, via $P(\hat{G}|X^{*})$, whereas $M$ is independent from it (i.e., $P(G,X|G^{*},X^{*})=P(G,X)$). The consideration that $\hat{G}$ is, in fact, resulting from a Bayesian procedure through a generative model $M$, instead of any—potentially nongenerative—model, will prove useful in the following sections.
From an information-theoretic perspective, data generation encodes information about the graph $G^{*}$ into potentially noisy observations $X^{*}$, while network reconstruction decodes these observations back into a graph $\hat{G}$ as shown in Fig. 1(a–b). The encoding of $G^{*}$ into $X^{*}$ is generally lossy, meaning that only a fraction of its information can be recovered; the rest being lost in the process. In turn, any reconstruction model $M$ different in distribution from $M^{*}$ may therefore recover a fraction of the reconstructible information while potentially introducing spurious information through their inductive biases [see Fig. 1(c–d)], resulting in a degradation of performance. This is well shown in Fig. 2 through reconstruction performance, where we compare the area under the receiver operating characteristic curve (AUC) of the TDG reconstruction model with the AUC of three heuristic reconstruction algorithms. In this example, the random graph of the TDG model is a simple Erdős-Rényi (ER) graph, whose graph prior is
@@
with $E$ being the (given) number of edges in the graph (see Appendix IV.1 for details). Then, the data consists in time series of binary node activities, generated using the Glauber dynamics model Glauber (1963) in panel (a) and the Susceptible-Infected-Susceptible (SIS) dynamics model Pastor-Satorras et al. (2015) in panel (b) (see Appendix IV.2 for details). As a comparison, we use three different well-known reconstruction algorithms: the correlation matrix method Kramer et al. (2009), Granger causality method Schreiber (2000) and the transfer entropy method Seth (2005) (see Appendix IV.3 for details). The results in Fig. 2 show quite unambiguously and unsurprisingly that the TDG model outperforms the reconstruction heuristics.
Yet, even the TDG reconstruction model cannot reconstruct the graph perfectly. For instance, in Fig 2(b), the AUC of the TDG model tends to $\frac{1}{2}$—equivalent to random guessing—when the coupling also goes to zero. In this scenario, $X^{*}$ and $G^{*}$ are independent and it is actually impossible to reconstruct the graph, since any graph could have generated the data with the exact same probability. The same phenomenon occurs to a lesser extent for the other coupling values as well as for the SIS dynamics, where the TDG model performance is imperfect for every infection probability. These imperfections are attributed to the lost information in the encoding of $G^{*}$; no model can extract more information than what is contained in $X$. In practice, the encoding’s loss stems from many sources, for example noise in the dynamics and degeneracy, where many networks lead to similar dynamics. The degeneracy phenomenon is well-established in computational neuroscience Prinz et al. (2004); Cropper et al. (2016) and has more recently appeared in network science Prasse and Mieghem (2022) too. A reconstruction limit independent of the reconstruction algorithm clearly exists, where a perfect reconstruction is simply not attainable even in the best-case scenario. This is a key insight that we will explore in the following sections (especially Sec. II.3).
II.1.2 Reconstructing a single edge
To gain better intuition about this reconstruction limit, we consider the reconstruction of a graph that may only contain a single edge. Let $G^{*}$ be a random graph of two nodes, that may be connected by a single edge with probability $p$, and disconnected with probability $1-p$. This edge is observed through a noisy process $X^{*}=(X_{1},...,X_{T})$ with $T$ time steps, where $X_{i}$ is a binary variable that takes the value $1$ if the edge has been observed and 00 otherwise. We assume that the noisy process can induce true positives and false positives, each with known probabilities $q$ and $r$, respectively, making the reconstruction problem more challenging. The likelihood $P(n|a)$ that the edge has been observed $n$ times, given that it is present ($a=1$) or not ($a=0$), is a binomial distribution:
@@ II.1.2 Reconstructing a single edge
Note that this model possesses a symmetry where interchanging $q$ and $r$ and mapping $a\to 1-a$ leaves the likelihood invariant. However, we avoid this non-identifiability issue by not inferring $p$ and $r$.
@@
Figure 3 shows the behavior of $P(a=1|n)$ when varying the number $n$ of times the edge is observed and the true positive probability $q$. Assuming that $r<q$, observed edges are mostly true positives and thus the edge is predicted to exist if $n$ is sufficiently large; otherwise, it is not since the expected number of true and false positives don’t match the observations. Conversely, if $r>q$, then most observed edges are false positives, meaning that $g^{*}$ is more likely to contain an edge when $n$ is small. Interestingly, the edge becomes more challenging to reconstruct as $q$ gets closer to $r$, where the probability to reconstruct the edge approaches $\frac{1}{2}$ (see Fig. 3(a)). In this regime, $a$ and $1-a$ are interchangeable and it becomes impossible to tell if the edge exists or not—any attempt at reconstructing this graph would be unfruitful. This is precisely the intuition we want to capture with the reconstruction limit: When is there enough information to properly reconstruct the structure, or to what extent is a system’s structure reconstructible? In the next section, we present an information-theoretic framework that quantifies this limit.
II.2 Information-theoretic reconstruction limits
As discussed above, we can think of the TDG process $X^{*}$ as a noisy encoding of the true graph $G^{*}$. This amount of encrypted information is fundamentally limiting our ability to reconstruct $G^{*}$ accurately; it is impossible to recover more information than what is contained in the data. This also means that the limit is independent of the reconstruction models or algorithms. Any reconstruction algorithm therefore aims to extract as much of the encrypted information as possible, some being more efficient than others.
II.2.1 Entropy